OpenStudy (anonymous):

Find the volume of a frustum of a pyramid with square base of side 12, square top of side 7 and height 12.

OpenStudy (anonymous):

you know this formula ?\[V _{F}=\frac{ 1 }{ 3 }h(A_{B}+A_{B^{'}}+\sqrt{A_{B}A_{B^{'}}})\]

OpenStudy (anonymous):

Can you work it out for me?

OpenStudy (anonymous):

Ab=12*12; Ab'=7*7 as well as h=12 Thus insert them in the formula

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

12*12 = 144 7*7 = 49

OpenStudy (anonymous):

?

OpenStudy (anonymous):

insert the above numbers in their respective places of the formula and solve it you must get this finally 1108